Grade 12Physical Chemistry

When molten LiCl is electrolyzed, lithium metal is formed at the cathode. If current efficiency is 75% then how many grams of lithium are liberated when 1930C of charge pass through the cell?

Hrishabh Singhal
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
Current efficiency = current passed actually * 100/ current passed experimently
wt metal = E " c * t /96500
Applied charge = 1930 C
Efficiency = 75%
Word chrged = 1930 * 75% = 1447.5 C
96500 C liberate 1 mole
1 C = 1/96500 moles
1447.5 C = 1447.5 /96500